Prepare 2 liters of a 4 molal solution of MgSO4 in water. The MgSO4 is the solute and water is the solvent. Use rounded numerical atomic weights instead of exact ones with decimal values.

a) how much MgSO4 should be weighed out?

b) how many molecules of MgSO4 should be weighed out?

c)How many individual atoms of solute are present?

d) how much water should be added?


1
Expert's answer
2021-09-20T06:01:32-0400

a) Molar mass of MgSO4 = 120.366 g/mol

2000/120.366 = 16.62/4 = 4.154moles

= 4.154× 120.366 = 500g

b) 4.154 × 6.023×1023 = 2.502 × 1024molecules